OpenStudy (anonymous):

Solve the equation. Be sure to show all of your work. -2x – 5 = -10

OpenStudy (anonymous):

multiply it by -1

OpenStudy (anonymous):

-2x-5 =-10 -2x-5+5=-10+5 -2x=-5 -2x*(-1)=-5*(-1) 2x=5 x=(5/2)

OpenStudy (anonymous):

-2x-5=-10 -2x=-10+5=-5 x=5/2 negative /negative=positive